# Beyond Chinchilla-Optimal: Accounting for Inference in Language Model Scaling Laws

### 31 Dec 2023 v1

@@ 4 Conclusion
In this work, we modify the Chinchilla scaling laws to account for both the computational and real-world costs of inference. As inference demand approaches pre-training data size, the additional cost pushes the optimal parameters-to-tokens ratio towards smaller and longer-trained models.
We make strong assumptions about the Chinchilla scaling laws and our analysis only applies insofar as these laws hold true. Further work is needed to experimentally validate our formulas and determine if scaling laws apply in the extreme ranges, where pre-training tokens exceed model parameters by orders of magnitudes.
@@ Acknowledgements
We thank Sasha Doubov for helpful discussions and Daya Khudia, Mihir Patel, and Linden Li for their feedback on the manuscript.
@@ Appendix A No Analytic Solution for Inference-Compute Optimality
Given a fixed model quality and inference demand, there exists no analytic solution for the compute-optimal model size and pre-training token count according to the third Chinchilla law, after accounting for the computational cost of inference.
@@
This constraint, from the third Chinchilla law, ensures we are minimizing compute while fixing model quality (pre-training loss). $A=406.4,B=410.7,E=1.69,\alpha=0.336$, and $\beta=0.283$ are constants determined empirically by Hoffmann et al. [2022]. [^2] The Chinchilla paper reports $\alpha=0.34$ and $\beta=0.28$. However, these are rounded values; to better fit the results reported in Table A.3 of Hoffmann et al. [2022], we use $\alpha=0.336$ and $\beta=0.283$, as in De Vries [2023].
@@
Thus, determining $D_{\text{tr}}$ as a function of $D_{\text{inf}}$ and $\ell$ involves finding the roots of equations of the form $ax^{-1.283}+756.6x^{-0.283}+c=0$ for arbitrary $a$ and $c>0$, which is not possible in general. ∎
@@ B.2 Cost-Optimal Results
We show additional results from our cost-optimality analysis in Sec. 3. In Table 2, we show the total training plus inference costs for Chinchilla models of various sizes at different levels of inference demands. We then calculate costs for equivalent-quality (i.e. same pre-training loss) cost-optimal models and show the overall savings. We use the same settings from Figure 2, designed to mimic a typical real-world deployment: training and inference input at 50% MFU, generation at 1% MFU Korthikanti et al. [2022], Pope et al. [2022]. Each inference request has 70 input tokens and 215 output tokens, in accordance with averages from the LMSYS-Chat dataset of 1M inference requests from Zheng et al. [2023]. Costs are calculated assuming training and inference on A100-80GB and A100-40GB accelerators, respectively. We further assume the model parameters are quantized to eight-bit integers prior to inference, which is commonly done with no quality reduction Xiao et al. [2023]. All costs are reported in US dollars.
@@ B.3 GPU Details
GPU pricing varies based on vendor and fluctuates over time. At the time of writing, an A100-40GB costs USD \$1.10/hr and an A100-80GB costs \$1.50/hr on Lambda Labs Labs [2023]. We use these values in our cost analysis in Sec. 3 and in Table 2. Both variants have a peak performance of $3.12\times 10^{14}$ dense FP16/BF16 operations and $6.24\times 10^{14}$ INT8 operations per second NVIDIA [2021].

### 18 Jul 2024 v2

@@ 8 Conclusion
In this work, we modify the Chinchilla scaling laws to account for both the computational and real-world costs of inference. As inference demand approaches pre-training data size, the additional cost pushes the optimal tokens-to-parameters ratio towards smaller and longer-trained models.
We experimentally validate the hypothesis that very small models, trained on enough data, can match larger Chinchilla-style ones. Both in terms of loss and downstream metrics, model quality improves as tokens per parameter increase. Further work is needed to show if this scales beyond 10,000 tokens per parameter, or at larger model sizes. Still, our results show that practitioners operating in inference-heavy regimes, or with limited deployment memory, can scale training duration considerably longer than current literature suggests and see quality improvements.
Finally, we show evidence that the Chinchilla parametric coefficient fitting procedure overestimates the reduction from additional training data when applied to extremely data-heavy training runs. More work is needed to develop scaling laws that apply precisely at a wide range of ratios.
@@ Acknowledgements
We thank Daya Khudia for his support and Mihir Patel and Linden Li for their feedback on this manuscript.
Impact Statement
This paper presents work whose goal is to reduce the costs of producing large language models. LLMs have the potential to produce factual, high-quality text, but also biased or harmful outputs. By reducing the costs train capable LLMs, we make them more accessible to scientific researchers, industry, and the general population.
@@ Appendix A No Analytic Solution for Inference-Compute Optimality
Given a fixed model quality and inference demand, there exists no general analytic solution for the compute-optimal model size and pre-training token count according to the third Chinchilla law, after accounting for the computational cost of inference.
@@
This constraint, from the third Chinchilla law, ensures we are minimizing compute while fixing model quality (pre-training loss). $A=406.4,B=410.7,E=1.69,\alpha=0.336$, and $\beta=0.283$ are constants determined empirically by Hoffmann et al. (2022). [^2] The Chinchilla paper reports $\alpha=0.34$ and $\beta=0.28$. However, these are rounded values; to better fit the results reported in Table A.3 of Hoffmann et al. (2022), we use $\alpha=0.336$ and $\beta=0.283$, as in De Vries (2023).
@@
Thus, determining $D_{\text{tr}}$ as a function of $D_{\text{inf}}$ and $\ell$ involves finding the roots of equations of the form $ax^{-1.283}+756.6x^{-0.283}+c=0$ for arbitrary $a$ and $c>0$, which is not possible in general.
The best-fit constants $A,B,E,\alpha$ and $\beta$ vary based on the exact dataset and model architecture. Outside of a handful of special-case values of $\beta$ for which Eq. 12 can be manipulated into a low-degree polynomial, it is intractable. ∎
@@ B.2 Cost-Optimal Results
We show additional results from our cost-optimality analysis in Sec. 6. In Table 3, we show the total training plus inference costs for Chinchilla models of various sizes at different levels of inference demands. We then calculate costs for equivalent-quality (i.e. same pre-training loss) cost-optimal models and show the overall savings. We use the same settings from Figure 6, designed to mimic a typical real-world deployment: training and inference input at 50% MFU, generation at 1% MFU (Korthikanti et al., 2022; Pope et al., 2022). Each inference request has 70 input tokens and 215 output tokens, in accordance with averages from the LMSYS-Chat dataset of 1M inference requests from Zheng et al. (2023). Costs are calculated assuming training and inference on A100-80GB and A100-40GB accelerators, respectively. We further assume the model parameters are quantized to eight-bit integers prior to inference, which is commonly done with no quality reduction (Xiao et al., 2023). All costs are reported in US dollars.
@@ B.3 GPU Details
GPU pricing varies based on vendor and fluctuates over time. At the time of writing, an A100-40GB costs USD \$1.10/hr and an A100-80GB costs \$1.50/hr on Lambda Labs (Labs, 2023). We use these values in our cost analysis in Sec. 6 and in Table 3. Both variants have a peak performance of $3.12\times 10^{14}$ dense FP16/BF16 operations and $6.24\times 10^{14}$ INT8 operations per second (NVIDIA, 2021).
Appendix C Model Training
We train MPT-style transformer models (MosaicML, 2023) ranging in size from 150M to 6B parameters, on token/parameter ratios from 10 to 10,000. In Table 4, we provide training configuration details. All models were trained with ALiBi (Press et al., 2022), Grouped Query Attention (Ainslie et al., 2023), the Lion optimizer (Chen et al., 2023) ($\beta_{1}=0.9,\beta_{2}=0.95$) with weight decay equal to the learning rate, cosine warmup ($\alpha_{f}=0.1$) with a duration equal to 3 times the number of model parameters, and norm gradient clipping (threshold = 1). A maximum sequence length of 4096 tokens was used. A smaller batch size was used for smaller models so that low-token-count training runs see enough update steps to learn properly. For all experimental results in this work, we use the smoothed final training loss over the last ten batches to reduce noise from minor batch-level variations.
In the rightmost column of Table 4, we list our experiment training token ratios. Note that for the larger models, our training durations are limited by our computational resources.
Table 4: Model Training Configurations.
Name Params d_model n_heads n_layers Learning Rate Batch Size Tokens/Parameter
150M 151M 768 12 12 4.603e-4 160333The 10,000 token/parameter 150M model was trained with batch size 960 due to compute issues. 10,15,20,30,50,75,100,
250,500,1000,5000,10000
370M 367M 1024 16 24 3.453e-4 320 10,15,20,30,50,75,100,
250,500,1000
750M 749M 1536 12 24 2.302e-4 480 10,15,20,30,50,75,100,250,500
1.3B 1.26B 2048 16 24 1.726e-4 960 10,15,20,30,50,75,100,250
2.5B 2.46B 2560 20 32 1.381e-4 960 10,15,20,50,100,250,500
6B 6.05B 4096 32 32 8.632e-5 960 20
Appendix D Experiment Results
We present more results from our experiments described in Section 3. In Figure 7, we show aggregate results for each Gauntlet category. Within each category, every task (the tasks are enumerated in Sec. 3) is weighted equally and averaged after subtracting out baseline and normalizing so the maximum achievable accuracy is 1. Note that for some categories (e.g. Symbolic Problem Solving), all models achieve nearly zero performance on all tasks, resulting in low correlation between accuracy and token/parameter ratios (see Barton (2024)).
Figure 7: Per-category Gauntlet results.